The trial balances shown below are before and after adjustment for Pharoah Company at the end of its fiscal year.

PHAROAH COMPANY
Trial Balance
August 31, 2017

Before
Adjustment

After
Adjustment

Dr. Cr. Dr. Cr.

Cash

$10,680 $10,680

Accounts Receivable

8,770 9,360

Supplies

2,990 1,910

Prepaid Insurance

4,210 2,960

Equipment

16,710 16,710

Accumulated Depreciation—Equipment

$3,760 $4,960

Accounts Payable

5,080 5,080

Salaries and Wages Payable

0 1,850

Unearned Rent Revenue

1,910 1,160

Common Stock

19,980 19,980

Retained Earnings

5,330 5,330

Dividends

2,800 2,800

Service Revenue

34,000 34,590

Rent Revenue

12,180 12,930

Salaries and Wages Expense

16,130 17,980

Supplies Expense

0 1,080

Rent Expense

19,950 19,950

Insurance Expense

0 1,250

Depreciation Expense

0

1,200

$82,240

$82,240

$85,880

$85,880


Prepare the adjusting entries that were made.
Answer #1

Solution

Pharoah Company

Adjusting Entries:

Date

Account Titles and Explanation

Debit

Credit

August 31, 2017

Accounts Receivable

$590

Sales Revenue

$590

(To record unbilled credit sales; 9,360 - 8,770 = 590)

August 31, 2017

Supplies Expense

$1,080

Supplies

$1,080

(To record cost of supplies used, 2,990 - 1,910 = 1,080)

August 31, 2017

Insurance Expense

$1,250

Prepaid Insurance

$1,250

(To record expired portion of prepaid insurance, 4,210 - 2,960 = 1,250)

August 31, 2017

Depreciation Expense - Equipment

$1,200

Accumulated Depreciation - Equipment

$1,200

(To record current year depreciation, 4,960 - 3,760 = 1,200)

August 31, 2017

Salaries and Wages Expense

$1,850

Salaries and Wages Payable

$1,850

(To record unpaid salaries)

August 31, 2017

Unearned Rent Revenue

$750

Rent Revenue

$750

(To record earned portion of unearned rent revenue, 1,910 - 1,160 = 750)